Solve the following system and present your answer below:
(5/2)x + y = 1/2
2x - (3/2)y = 5

Answer by Alan3354(69443)

(5/2)x + y = 1/2
2x - (3/2)y = 5
------------------
Eqns times 2:
5x + 2y = 1 Eqn A
4x - 3y = 10 Eqn B
------
15x+ 6y = 3 Eqn A times 3
8x - 6y = 20 Eqn B times 2
------------------------------- Add
23x = 23
x = 1
============
Sub for x in any eqn and solve for y